# colorlinks = true, linkcolor = blue, citecolor = blue, urlcolor = blue

### v1

@@ IV-A Robot Setup
For our experiment, we apply a UR5e robotic arm, which is equipped with an OnRobot RG2 gripper. An Intel RealSense D435 RGB-D camera is attached to the wrist of the UR5e robotic arm to obtain the RGB image and depth information. The robot is controlled via the Robot Operating System (ROS). In particular, we use the RRT-Connect algorithm for path planning.
@@ IV-B Outer Knowledge of Successful Rearrangement
It is important to note that this outer knowledge of successful rearrangements is not used for training the model but rather serves as a reference to enhance the LLM’s spatial reasoning capabilities. The external knowledge from a successful experiment includes both spatial and semantic information, such as the name of each object, the object’s center point, its bounding box, rotation, and the corresponding instructions. 10 arrangements were created manually using representations of kitchen items. Specifically, the image of the successful human-arranged object rearrangement is captured from a top-down view. Based on the objects shown in the image, we use SAM to segment the corresponding masks using location prompts. From these masks, we derive the centre point, bounding box, and rotation of each object. Seven of the scenarios contain two objects, while the remaining three contain multiple objects. Some examples are shown in the Fig 4. Our framework also has the potential to expand external knowledge by recording the spatial information of objects during successful rearrangements performed by robots. However, in our experiments, we use only these 10 arrangements as outer knowledge to highlight how they improve the results.
@@ IV-C User Study on Object Rearrangement
In our experiment, 15 arrangements are performed on 3 real-world scenes. The first scene consists of only an eggplant and a plate, with the robot manipulating a single object. The instructions for this scene are: (1) “put the eggplant on the right of the plate,” (2) “put the eggplant on the left of the plate,” (3) “put the eggplant in front of the plate,” (4) “put the eggplant behind the plate,” (5) “put the eggplant far away from the plate.” The second scene includes repetitive objects, i.e., two potatoes, to assess whether our method can effectively rearrange identical items and multiple objects. The instructions for the second scene are: (1) “put the potatoes on the plate,” (2) “put the potatoes beside the plate,” (3) “put one potato to the left of the plate and the other to the right,” (4) “put the potatoes far away from the plate,” (5) “put the potatoes together.”
@@
In the third scene, our objective is to determine if our method can effectively rearrange objects in a long sequential order. This scene includes several objects, such as a carrot, a potato, a pineapple, an eggplant and a plate. The instructions for the third scene are: (1) “put the eggplant on the plate, then beside the plate,” (2) “put eggplant beside the plate, then beside the carrot,” (3) “put the eggplant beside the potato, then put the eggplant on the plate,” (4) “put the eggplant beside the carrot, then far away from the carrot,” (5) “put the eggplant on the right of the potato, then on the left of the pineapple.” In particular, we use the ChatGPT-4 [27] for spatial reasoning.
@@
For the evaluation of object rearrangement, our objective is to measure if the final arrangement of objects matches the given instructions. Since our framework aims to mimic the human thinking method, a straightforward method is to have a user study to gather feedback from humans, following the approach in [16] [8]. Specifically, 10 participants were invited to evaluate whether the final real-world scene created by the robot meets the given instructions. Two of the participants are female, while the others are male, with ages ranging from 25 to 33. Participants rated the results on a scale from 1 to 10, where 1 indicates “not acceptable” and 10 indicates “highly acceptable”. All users were presented with the same set of images and the set-up for every method was similar. The participants evaluated the results from 8 different methods, providing a total of 1200 ratings.
@@
As shown in Table I, we present a comparison of our results with other baseline methods. Since most existing approaches rely on large training datasets, they are not directly comparable to our method in a zero-shot manner. To address this, two heuristic-based baselines are selected for comparison. In the Random baseline, the robot picks objects and places them at random locations. In the Geometric baseline, the robot arranges the relevant objects in a horizontal line with a certain gap, ensuring their centre points are aligned at the same level. Moreover, we compare the results with an ablated structure that does not use the reference for spatial reasoning, which follows a similar structure as described in [21]. In [21], the LLM is used for categorisation, whereas we employ the LLM for spatial reasoning.
As shown in Table I, our proposed method achieves the highest average evaluation score of 9.09, outperforming other methods. This suggests that the robot can rearrange objects in a manner that closely resembles human-like behaviour. Furthermore, the results of our method without reference outperform those of heuristic-based baselines, highlighting the strong spatial reasoning capabilities of the LLM. Based on the LLM, the mean value increases by 1.16 while using the reference during spatial reasoning, indicating that our framework effectively transfers knowledge from prior successful experiences to facilitate more efficient object rearrangement. The examples in Fig. 5 also illustrate that the use of reference enhances the accuracy of object placement and enables a more reasonable gap between objects. However, the performance for single object rearrangement is lower than for other tasks. This is likely because the instructions include terms like “in front of” and “behind”, which make it more challenging for the LLM to accurately infer object positions. More results are available on our webpage at: https://sites.google.com/view/learnpast
IV-D Comparison between Different Language Models
In this experiment, we apply various language models to evaluate how their capabilities affect the prediction of goal positions in object rearrangement. Instead of using ChatGPT-4, we employ smaller LLMs to determine whether our proposed method can enhance performance even with models that require lower computational resources, thereby making their results comparable to those of more computationally intensive models.
Specifically, we use different backbone models Mistral-7B [28] and Llama3-8B [29] for spatial reasoning. These models are applied in both configurations, with and without reference, respectively. From Table II, we can observe that using a reference improves performance across all models in most tasks. In particular, the mistral model gains significant improvement from references in tasks involving multiple object rearrangement, with an increase of 3.64 in the score. However, Mistral and Llama3 show more variability with a higher standard deviation, which indicates that they are less consistent compared to ChatGPT-4. It can be seen ChatGPT-4 outperforms smaller models in almost every scenario, both with and without reference, but the performance gap narrows when reference is provided, compared to the other two models. Given that ChatGPT4 is significantly larger than Mistral-7B and Llama3-8B, the result suggests that the use of references can mitigate the limitations caused by decreased computational capacity.
@@ IV-D Comparison between Different Language Models
Random 1.82±$\pm{1.61}$ 5.48±$\pm{2.50}$ 4.84±$\pm{3.40}$ 4.05
Geometric 2.98±$\pm{3.41}$ 4.76±$\pm{3.41}$ 4.80±$\pm{4.01}$ 4.18
w/o reference [21] 5.38±$\pm{3.74}$ 9.02±$\pm{1.82}$ 9.4±$\pm{1.31}$ 7.93
Ours 8.58±$\pm{\textbf{2.20}}$ 9.14±$\pm{\textbf{1.65}}$ 9.54±$\pm{\textbf{0.86}}$ 9.09
@@
mistral w/o reference 3.58±$\pm{3.39}$ 3.52±$\pm{3.01}$ 7.78±$\pm{2.90}$ 4.96
llama3 w/o reference 5.36±$\pm{3.36}$ 6.00±$\pm{4.21}$ 6.92±$\pm{3.77}$ 6.09
ChatGPT4 w/o reference 5.38±$\pm{3.74}$ 9.02±$\pm{1.82}$ 9.4±$\pm{1.31}$ 7.93
mistral w/ reference 5.02±$\pm{3.62}$ 7.16±$\pm{3.24}$ 9.14±$\pm{1.43}$ 7.11
llama3 w/ reference 7.74±$\pm{3.22}$ 6.74±$\pm{2.83}$ 8.56±$\pm{2.12}$ 7.68
chatGPT4 w/ reference 8.58±$\pm{\textbf{2.20}}$ 9.14±$\pm{\textbf{1.65}}$ 9.54±$\pm{\textbf{0.86}}$ 9.09
@@ V Conclusion
In this paper, we propose an LLM-based framework for language-conditioned object rearrangement. By using past successful rearrangements as a reference, the robot can more accurately predict target positions for object placement, thereby mimicking the human reasoning process (see [4][5]). The results demonstrate that the use of references improve performance across multiple dimensions, including the rearrangement of a single object, multiple objects and tasks involving instructions with a sequential order. Additionally, the application of outer knowledge can help overcome the limitations caused by computational constraints, making it easier for practical applications in local deployment.
Nevertheless, the proposed method has some limitations that may affect its performance in real-world applications. Firstly, the scenario we designed operates primarily on a 2D surface, rather than in a 3D space, which limits its applicability in the physical world. Additionally, due to the spatial reasoning constraints of LLMs, we focus only on inferring the goal position but lack predicting the goal orientation for the arrangement. The objects used in the space are relatively sparse and this could pose a challenge in more clustered environments. Future improvements would involve extending the method to 3D rearrangement and addressing object rearrangement in clustered settings.

### v2

@@ IV-A Robot Setup
For our experiment, we apply a UR5e robotic arm, which is equipped with an OnRobot RG2 gripper. An Intel RealSense D435 RGB-D camera is attached to the wrist of the UR5e robotic arm to obtain the RGB image and depth information. The robot is controlled via the Robot Operating System (ROS) and we use the RRT-Connect algorithm for path planning.
@@ IV-B Outer Knowledge of Successful Rearrangement
It is important to note that this outer knowledge of successful rearrangements is not used for training the model but rather serves as a reference to enhance the LLM’s spatial reasoning capabilities. The external knowledge from a successful experiment includes both spatial and semantic information, such as the name of each object, the object’s centroid, its bounding box, rotation, and the corresponding instructions. 10 arrangements were created manually using representations of kitchen items. Specifically, the image of the successful human-arranged object rearrangement is captured from a top-down view. Based on the objects shown in the image, we use SAM to segment the corresponding masks using location prompts. From these masks, we derive the centroid, bounding box, and rotation of each object. Eight of the scenarios contain two objects, while the remaining two contain multiple objects. Some examples are shown in the Fig 5. Our framework also has the potential to expand external knowledge by recording the spatial information of objects during successful rearrangements performed by robots. This can then be used as successful past experiences for future tasks similarly to Fig. 2. However, in our experiments, we use only these 10 arrangements as outer knowledge to highlight how they improve the results. More details are available on our webpage at: https://sites.google.com/view/learnpast
@@ IV-C User Study on Object Rearrangement
In our experiment, 15 arrangements are performed on 3 real-world task scenarios. The first scene consists of only an eggplant and a plate, with the robot manipulating a single object. The instructions for this scene are: (1) “put the eggplant on the right of the plate,” (2) “put the eggplant on the left of the plate,” (3) “put the eggplant in front of the plate,” (4) “put the eggplant behind the plate,” and (5) “put the eggplant far away from the plate.” The second scenario include repetitive objects, i.e., two potatoes, to assess whether our method can effectively rearrange identical items and multiple objects. The instructions for the second scene are: (1) “put the potatoes on the plate,” (2) “put the potatoes beside the plate,” (3) “put one potato to the left of the plate and the other to the right,” (4) “put the potatoes far away from the plate,” (5) “put the potatoes together.”
@@
In the third scene, our objective is to determine if our method can effectively rearrange objects in a long sequential order. This scene includes several objects, such as a carrot, a potato, a pineapple, an eggplant and a plate. The instructions for the third scene are: (1) “put the eggplant on the plate, then beside the plate,” (2) “put eggplant beside the plate, then beside the carrot,” (3) “put the eggplant beside the potato, then put the eggplant on the plate,” (4) “put the eggplant beside the carrot, then far away from the carrot,” (5) “put the eggplant on the right of the potato, then on the left of the pineapple.”
@@
For the evaluation of object rearrangement, our objective is to measure if the final arrangement of objects matches the given instructions. Since our framework aims to mimic human thinking and ensure that the rearrangement aligns with human preferences, a straightforward approach is to conduct a user study to gather human feedback, following the methodology in [13] and [7]. Specifically, 15 participants were invited to evaluate whether the final real-world scene created by the robot meets the given instructions. Four of the participants were female, while the others were male, with ages ranging from 25 to 33. Participants rated the results on a scale from 1 to 10, where 1 indicates “not acceptable” and 10 indicates “highly acceptable”. All users were presented with the same set of images and the set-up for every method was similar. The participants evaluated the results from 9 different methods (shown in Table I and Table II), providing a total of 2025 ratings.
@@
As shown in Table I, we present a comparison of our results with other baseline methods. Since most existing approaches rely on large training datasets, they are not directly comparable to our method in a zero-shot manner. To address this, two heuristic-based baselines are selected for comparison. In the Random baseline, the robot picks objects and places them at random locations. In the Geometric baseline, the robot arranges the relevant objects in a horizontal line with a certain gap, ensuring their centre points are aligned at the same level. We also adapted Dream2Real [31] to fit our tasks and use it as a baseline. Dream2Real uses sampling to place objects in various locations in order to create a scene aligned with the goal caption. Additionally, we compare our results with an ablated structure that does not use referencing or templating for spatial reasoning. This follows a similar structure described in [23] where the LLM is used for categorisation. However, we applied the LLM for spatial reasoning in this work.
As illustrated in Table I, our proposed method achieves the highest average evaluation score of 9.14, outperforming all other methods. The results also demonstrate superiority across all three evaluation scenarios: single-object placement, multi-object arrangement, and sequential ordering, indicating that the robot can effectively reason about spatial relationships and arrange objects in a reasonable place. Notably, even when operating without reference guidance, our framework maintains competitive performance (7.93 mean score), outperforming the sampling-based approach [31] and heuristic-based baselines, which demonstrates the inherent effectiveness of our LLM-driven architecture. By incorporating references from successful experiences, our approach further enhances spatial accuracy, increasing the mean evaluation score by 1.21 compared to results without reference. This suggests that our approach effectively leverages prior successful experiences to enhance knowledge transfer, facilitating more efficient and intelligent object rearrangement. The results in Fig. 6 also illustrate that the use of reference enhances the accuracy of object placement and enables a more reasonable gap between objects. However, it is observed that the performance in single-object rearrangement is lower compared to multiple objects and sequential order tasks. This is likely because the instructions for single-object placement contain more precise spatial terms such as “in front of”, “behind”, “left of”, and “right of”, which pose a greater challenge for the LLM in accurately inferring object positions. In contrast, the other two tasks involve less precise spatial terms, such as “beside” and “far away from”, which are easier for the model to perform reasoning.
@@
Random 1.63±$\pm{1.39}$ 5.57±$\pm{2.26}$ 4.61±$\pm{3.44}$ 3.94
Geometric 2.92±$\pm{3.48}$ 4.65±$\pm{3.19}$ $\pm{4.08}$ 4.12
Dream2 Real [31] 6.16±$\pm{3.24}$ 6.95±$\pm{2.27}$ 7.33±$\pm{2.77}$ 6.81
Ours w/o reference [23] 5.51±$\pm{3.71}$ 8.92±$\pm{1.85}$ 9.37±$\pm{1.17}$ 7.93
Ours 8.67±$\pm{\textbf{2.18}}$ 9.24±$\pm{\textbf{1.44}}$ 9.51±$\pm{\textbf{0.86}}$ 9.14
IV-D Comparison between Different Language Models
In this experiment, we apply various language models to evaluate how their capabilities affect the prediction of goal positions in object rearrangement. Instead of using an “extra” Large Language Model like ChatGPT-4, we used smaller LLMs to assess whether our proposed framework can still achieve adequate performance with models that require lower computational resources and with results comparable to more resource-intensive models.
Specifically, we use different backbone models Mistral-7B [32] and Llama3-8B [33] for spatial reasoning. These models are applied in two configurations: with and without reference. From Table II, we can observe that using a reference improves performance across all models in most tasks, with an increase of 1.21, 1.52 and 2.24 for ChatGPT4, Llamma3 and Mistral respectively. In particular, the mistral model gains significant improvement from references in tasks involving multiple object rearrangement, with an increase of 3.78 in the score. It can be seen that ChatGPT-4 outperforms the smaller models across nearly all scenarios, regardless of whether a reference is used. However, the performance gap between ChatGPT-4 and the smaller models narrows when a reference is provided. Given that ChatGPT4 is significantly larger than Mistral-7B and Llama3-8B, the result suggests that the use of references can mitigate the limitations caused by decreased computational capacity.
Additionally, to make the results more clearer, we present the success rate of rearrangement in Fig.7. We define a rearrangement as successful if its evaluation score is greater or equal to 7 and calculate the overall success rate as the number of ratings $\geq 7$ divided by the total number of ratings. The results follow a similar trend where the ChatGPT using reference achieves the highest success rate of 95.11%. Additionally, it can be seen that incorporating reference guidance improves rearrangement performance by 18.22%, 18.67% and 29.34% for ChatGPT-4, Llama3, and Mistral respectively. This demonstrates that reference is particularly beneficial for models with lower computational requirements.
@@ IV-D Comparison between Different Language Models
Mistral w/o reference 3.69±$\pm{3.42}$ 3.21±$\pm{2.75}$ 7.69±$\pm{2.94}$ 4.87
Llama3 w/o reference 5.39±$\pm{3.34}$ 6.00±$\pm{4.19}$ 6.84±$\pm{3.86}$ 6.08
ChatGPT4 w/o reference 5.51±$\pm{3.71}$ 8.92±$\pm{1.85}$ 9.37±$\pm{1.17}$ 7.93
Mistral w/ reference 5.24±$\pm{3.70}$ 6.99±$\pm{3.37}$ 9.09±$\pm{1.32}$ 7.11
Llama3 w/ reference 7.83±$\pm{3.21}$ 6.48±$\pm{2.93}$ 8.48±$\pm{2.11}$ 7.60
ChatGPT4 w/ reference 8.67±$\pm{\textbf{2.18}}$ 9.24±$\pm{\textbf{1.44}}$ 9.51±$\pm{\textbf{0.86}}$ 9.14
Figure 7: Overall success rate. Success rates using different LLMs. The blue and orange bars represent the results with and without using references, respectively.
@@ V Conclusion
In order to achieve flexibility and efficiency, humans make use of past experiences for reasoning in novel task situations. On the other hand, most of the current methods in robotics make use of large datasets to train robots to achieve required goal states from a starting state, which reduces the flexibility and efficiency of robots in dealing with varying situations. To bridge this gap, this paper proposes a framework that leverages past successful rearrangements as references or templates to achieve efficient and flexible object manipulation, thereby mimicking human reasoning (see [9][10]). This results in better prediction of target positions for accurate object placement. Our results demonstrate that the use of references improve performance across multiple dimensions, including the rearrangement of a single object, multiple objects and tasks involving instructions with a sequential order. Additionally, the application of outer knowledge can help overcome the limitations caused by computational constraints, making it attractive for local deployment in practical applications. Nevertheless, the proposed method has some limitations that may affect its performance in real-world applications. Firstly, the scenario we designed operates primarily on a 2D surface, rather than in a 3D space, which limits its applicability in the physical world. Additionally, the objects used in the space are relatively sparse and this could pose a challenge in more clustered environments. Future improvements would involve extending the method to 3D rearrangement and addressing object rearrangement in clustered settings.